# Multi-UAV Path Following using Vector-Field Guidance

### 20 Apr 2026 v1

@@ Abstract
This paper presents a decentralized, collision-free framework for path following guidance of multiple uncrewed aerial vehicles (UAVs), while maintaining uniform spacing along a reference path. A vector field-based guidance law is employed to drive each UAV toward the reference path. A rotational repulsion mechanism, utilizing relative distance and bearing between UAVs, is proposed to avoid collisions during convergence to the path, and an inter-UAV spacing error-based velocity control law is presented to achieve uniform separation along the path. Analytical guarantees are established for collision avoidance and convergence of the inter-UAV spacing errors to zero, ensuring uniform separation along the path. Numerical simulations demonstrate the efficacy of the proposed method.
@@ 1 Introduction
With the rapid deployment of uncrewed aerial vehicles (UAVs) in civilian and industrial applications, such as aerial surveillance, package delivery, and infrastructure inspection, there has emerged a need for safe and robust trajectory planning methods. In a structured common airspace, UAVs are often required to fly along predefined routes, necessitating the design of collision avoidance strategies. In such scenarios, maintaining a uniform inter-UAV separation along the desired path is equally important, as it facilitates deterministic UAV traffic flow, efficient airspace utilization, and simplified conflict-resolution methods at intersections (Nagrare et al. (2024); He et al. (2025)).
@@
To address multi-UAV path-following along a designated non-closed path, various formation control strategies have been proposed in the literature. In Ghommam and Saad (2009), a Serret-Frenet formulation-based path-following control law coupled with a speed adjustment mechanism is developed for the UAVs tracking a common reference in a synchronized manner. That approach is centralized, which may limit its applicability to a large swarm. A hybrid control strategy is proposed in Chen et al. (2021), where the path-following error for an agent is shown to continuously decrease within its coordination set defined by the agent’s dynamics and motion constraints. The approaches in Ghommam and Saad (2009); Chen et al. (2021) lack explicit treatment of inter-UAV collision avoidance guarantees. In Prodan et al. (2010); Chen et al. (2025), coordinated control strategies for following continuous curved paths ensure collision avoidance by incorporating inter-UAV distance constraints into Model Predictive Control frameworks. These approaches require solving high-dimensional optimization problems at each control step, thereby restricting utility for a large UAV swarm due to high computational burden. While the approach in Zuo et al. (2021) achieves the tracking of distinct parameterized paths for different agents and maintains temporal coordination between agents, it does not address convergence to a common path with uniform inter-agent spacing.
Vector field-based methods have been widely employed in multi-UAV systems for various applications, including traversal in obstacle-laden environments (Hegde and Panagou (2016); Pothen and Ratnoo (2017)), evader capture (Viet et al. (2012)), and patrolling missions (Gao et al. (2018)), owing to their advantages of being computationally light and robust to external disturbances. The guiding vector field-based path following method in Yao et al. (2021) ensures path-following along a reference closed path for multiple robots but lacks explicit inter-UAV collision avoidance. Motivated by the need for scalable and collision-free path-following of a common path by multiple UAVs, this work presents a decentralized coordination strategy that enables multiple UAVs to converge to common open paths by leveraging an arcsine-based vector-field guidance law, while maintaining safe inter-UAV separation and uniform spacing along the path. The contributions of this work are as follows:
@@ 2 Problem Statement
    the UAVs converge to the reference path $\mathcal{P}_{k}\;(k=\{1,2\})$,
    $\lim_{t\to\infty}\text{dist}(\mathbf{x}_{i}(t),\mathcal{P}_{k})=0$ (4)
    where $\text{dist}(\mathbf{x}_{i},\mathcal{P}_{k})$ denotes the minimum Euclidean distance between the $i$th UAV and $\mathcal{P}_{k}$.
@@
Further, following assumptions have been considered in this work:
@@ 3.2 Sinusoidal Path Following
where $k_{p}>0$ is the heading control gain.
@@ 4.1 Collision Avoidance Mechanism
    The term $\left(\dfrac{1}{d_{i,j}}-\dfrac{1}{R_{s}}\right)$ is a distance-dependent weigthing factor that controls the repulsive turning rate based on the proximity between the $i$th and $j$th UAVs. At $d_{i,j}=R_{s}$, $\omega_{i}^{\mathrm{rep}}=0$, and grows monotonically as $d_{i,j}$ decreases below $R_{s}$, generating an angular deviation away from the neighboring agent. This inverse-distance dependence is analogous to the gradient of classical potential-field repulsion functions Rimon (1990); Khatib (1986).
@@
The following theorem establishes the collision avoidance property of the repulsive angular component in (18) under the equal-speed assumption $v_{i}=v_{j}=v$. Additionally, the repulsion component is assumed sufficiently large to dominate the total angular velocity input in the case of a potential collision.
@@
For $k_{r}=k_{r}^{*}$, $\ddot{d}_{i,j}=0$ occurs at ${d}_{i,j}^{*}=d_{\mathrm{safe}}$. Therefore, a sufficiently large gain $k_{r}>k_{r}^{*}$ exists such that $d_{i,j}^{*}>d_{\mathrm{safe}}$, which drives the range rate $\dot{d}_{i,j}$ to positive values, before $d_{i,j}=d_{\mathrm{safe}}$.
@@ 4.2 Equispacing Velocity Control
where $d_{\mathrm{eq}}$ is the desired inter-UAV spacing and the arc-length path parameter $s$ is defined as follows:
@@
    For the straight path $\mathcal{P}_{1}:x=0$,
@@
where $\kappa\in(0,\,v_{\mathrm{nom}})$ is a spacing gain and $v_{\mathrm{nom}}$ is the nominal speed common to all UAVs, and the saturation structure ensures
@@
where $\Delta_{1}=0$ indicates that the first UAV’s speed is $v_{\mathrm{nom}}$. Since $\tanh(\cdot)$ is monotonically increasing, each term $\Delta_{i}\bigl(\tanh(\Delta_{i})-\tanh(\Delta_{i-1})\bigr)$ is non-negative, which gives $\dot{V}\leq 0$. By LaSalle’s invariance principle, all trajectories converge to the largest invariant subset of $\{\dot{V}=0\}$. Since each term in (36) is non-negative, $\dot{V}=0$ implies
@@
The path-parameter rate $\dot{s}_{i}$ equals the component of velocity tangent to the path, that is, $\dot{s}_{i}=v_{i}\cos(\psi_{i}-\chi_{i}^{p})$, which reduces to $v_{i}$ when $\mathbf{x}_{i}\in\mathcal{P}_{k}\;(k=\{1,2\})$. Therefore, the asymptotic convergence to the path established in Section 3 justifies the assumption $\dot{s}_{i}=v_{i}$ used in Proposition 2.
@@ 5 Simulation Results
In this section, the proposed method is validated through numerical simulations in MATLAB for straight-path and sinusoidal-curve following. The initial UAV positions are uniformly sampled within a rectangular region bounded by the lines y=20 m,y=20 m,x=20 m,x=20 $\text{ m},y=20\text${ $\text{ m},x=-20\text${ m}. The safety distance of the UAVs and activation radius in (18) are considered to be $d_{\text{safe}}=0.4$ m and $R_{s}=1.5$ m, respectively. The nominal speed $v_{\text{nom}}=3$ m/s and $\kappa=1$ m/s, and the gain values $k_{r}=3.6,k_{g}=0.05,k_{\psi}=2.3$. The desired uniform spacing $d_{eq}$ is considered as 4 m. Further, we define the inter-UAV distance parameter $\mathcal{E}$ as the minimum distance between any UAV pair at a given time, that is, $\mathcal{E}=\min_{i\neq j}d_{i,j}$.
Figure 2: UAV trajectories for straight path following.
Figure 3: UAV control input and spacing error for straight path following: angular velocity input (top), linear velocity (middle), and spacing error (bottom).
Figure 4: Inter-UAV distance parameter $\mathcal{E}$ for straight path and sinusoidal curve following.
For the first scenario, straight line path-following is considered $(\mathcal{P}_{1}:x=0)$. The UAV trajectories are depicted in Fig. 2, which shows UAV positions at $t=0,20$ s and $40$ s. Fig. 3 shows the angular and linear velocity profiles of different UAVs, and the spacing error between consecutive UAVs. The angular velocity for all UAVs converge to 0, while the linear velocity converges to $v_{\text{nom}}=3$ m/s, while remaining within the bounds $(v_{\text{nom}}-\kappa,v_{\text{nom}}+\kappa)=(2,4)$ m/s. Since $\mathcal{E}>d_{\mathrm{safe}}=0.4$ m at all times, as shown in Fig. 4, no collision occurs between UAVs.
@@ 6 Conclusion
This paper presents a collision avoidance strategy for multiple UAVs following an open reference path while maintaining equal spacing along it. Two types of reference paths, straight line and sinusoidal curve, are considered, and a vector-field guidance law based on an arcsine function of the cross-track error drives each UAV to its reference path. Collision avoidance is achieved using a rotational repulsion mechanism that drives each UAV away from its neighbors based on relative bearing and proximity. The uniform spacing between UAVs along the path is maintained using a linear velocity control law that regulates each UAV’s speed based on the spacing error. Numerical simulations with 15 UAVs are performed to show the effectiveness of the proposed method. Future work could explore different paths, such as closed curves and more complex paths, and extend path-following to include collision-avoidance analysis in three-dimensional scenarios.

### 21 Apr 2026 v2

@@ Abstract
This paper presents a decentralized, collision-free framework for path following guidance of multiple uncrewed aerial vehicles (UAVs), while maintaining uniform spacing along a reference path. A vector field-based guidance law is employed to drive each UAV toward the reference path. A rotational repulsion mechanism, utilizing relative distance and bearing between UAVs, is proposed to avoid collisions during convergence to the path, and an inter-UAV spacing error-based velocity control law is presented to achieve uniform separation along the path. Analytical guarantees are established for collision avoidance and for the convergence of the inter-UAV spacing errors to zero, ensuring uniform separation along the path. Numerical simulations demonstrate the efficacy of the proposed method.
@@ 1 Introduction
With the rapid deployment of uncrewed aerial vehicles (UAVs) in civilian and industrial applications, such as aerial surveillance, package delivery, and infrastructure inspection, there has emerged a need for safe and robust trajectory planning methods. In a structured common airspace, UAVs are often required to fly along predefined routes, necessitating the design of collision avoidance strategies. In such scenarios, maintaining a uniform inter-UAV separation along the desired path is equally important, as it facilitates deterministic UAV traffic flow, efficient airspace utilization, and simplified conflict resolution at intersections (Nagrare et al. (2024); He et al. (2025)).
@@
To address multi-UAV path-following along a designated non-closed path, various formation control strategies have been proposed in the literature. In Ghommam and Saad (2009), a Serret-Frenet formulation-based path-following control law coupled with a speed adjustment mechanism is developed for UAVs tracking a common reference in a synchronized manner. That approach is centralized, which may limit its applicability to large swarms. A hybrid control strategy is proposed in Chen et al. (2021), in which the path-following error for an agent is shown to continuously decrease within its coordination set, defined by the agent’s dynamics and motion constraints. The approaches in Ghommam and Saad (2009); Chen et al. (2021) lack explicit treatment of inter-UAV collision avoidance guarantees. In Prodan et al. (2010); Chen et al. (2025), coordinated control strategies for following continuous curved paths ensure collision avoidance by incorporating inter-UAV distance constraints into Model Predictive Control frameworks. These approaches require solving high-dimensional optimization problems at each control step, thereby restricting utility for a large UAV swarm due to the high computational burden. While the approach in Zuo et al. (2021) achieves the tracking of distinct parameterized paths for different agents and maintains temporal coordination between agents, it does not address convergence to a common path with uniform inter-agent spacing.
Vector field-based methods have been widely employed in multi-UAV systems for various applications, including traversal in obstacle-laden environments (Hegde and Panagou (2016); Pothen and Ratnoo (2017)), evader capture (Viet et al. (2012)), and patrolling missions (Gao et al. (2018)), owing to their advantages of being computationally light and robust to external disturbances. The guiding vector field-based path following method in Yao et al. (2021) ensures path-following along a reference closed path for multiple robots but lacks explicit inter-UAV collision avoidance. Motivated by the need for scalable and collision-free path-following of a common path by multiple UAVs, this work presents a decentralized coordination strategy that enables multiple UAVs to converge to a common open path by leveraging an arcsine-based vector-field guidance law, while maintaining safe inter-UAV separation and uniform spacing along the path. The contributions of this work are as follows:
@@ 2 Problem Statement
    the UAVs converge to the reference path $\mathcal{P}_{m}\;(m=\{1,2\})$,
    $\lim_{t\to\infty}\text{dist}(\mathbf{x}_{i}(t),\mathcal{P}_{m})=0$ (4)
    where $\text{dist}(\mathbf{x}_{i},\mathcal{P}_{m})$ denotes the minimum Euclidean distance between the $i$th UAV and $\mathcal{P}_{m}$.
@@
Further, the following assumptions have been considered in this work:
@@ 3.2 Sinusoidal Path Following
where $k_{\psi}>0$ is the heading control gain.
@@ 4.1 Collision Avoidance Mechanism
    The term $\left(\dfrac{1}{d_{i,j}}-\dfrac{1}{R_{s}}\right)$ is a distance-dependent weighting factor that controls the repulsive turning rate based on the proximity between the $i$th and $j$th UAVs. At $d_{i,j}=R_{s}$, $\omega_{i}^{\mathrm{rep}}=0$, and grows monotonically as $d_{i,j}$ decreases below $R_{s}$, generating an angular deviation away from the neighboring agent. This inverse-distance dependence is analogous to the gradient of classical potential-field repulsion functions (Rimon and Koditschek (1992); Khatib (1986)).
@@
The following proposition establishes the collision avoidance property of the repulsive angular component in (18) under the equal-speed assumption $v_{i}=v_{j}=v$. Additionally, the repulsion component is assumed sufficiently large to dominate the total angular velocity input in the case of a potential collision.
@@
For $k_{r}=k_{r}^{*}$, $\ddot{d}_{i,j}=0$ occurs at ${d}_{i,j}^{*}=d_{\mathrm{safe}}$. Since $\frac{(\sin\phi_{j}-\sin\phi_{i})^{2}}{(\sin^{2}\phi_{i}+\sin^{2}\phi_{j})}\leq 2$, a sufficient condition for $d_{i,j}^{*}>d_{\mathrm{safe}}$ for all geometries $(\phi_{i},\phi_{j})$ is $k_{r}>\frac{2v\,R_{s}}{R_{s}-d_{\mathrm{safe}}}$. Therefore, a sufficiently large gain $k_{r}$ exists such that $d_{i,j}^{*}>d_{\mathrm{safe}}$, which drives the range rate $\dot{d}_{i,j}$ to positive values, before $d_{i,j}=d_{\mathrm{safe}}$.
@@ 4.2 Equispacing Velocity Control
where $d_{\mathrm{eq}}$ and $s$ are the desired inter-UAV spacing and the path parameter, respectively. This predecessor-following spacing error follows the string topology studied in Swaroop and Hedrick (1996). In this work, the path parameter $s$ is defined as follows:
@@
    For the straight line path $\mathcal{P}_{1}:x=0$,
@@
where $\kappa\in(0,\,v_{\mathrm{nom}})$ is a spacing gain and $v_{\mathrm{nom}}$ is the nominal speed common to all UAVs, and the hyperbolic tangent function structure ensures
@@
where $\Delta_{1}=0$ indicates that the first UAV’s speed is $v_{\mathrm{nom}}$. Since $\tanh(\cdot)$ is monotonically increasing, each term $\Delta_{i}\bigl(\tanh(\Delta_{i})-\tanh(\Delta_{i-1})\bigr)$ is non-negative, which gives $\dot{V}\leq 0$. By LaSalle’s invariance principle (Khalil and Grizzle (2002)), all trajectories converge to the largest invariant subset of $\{\dot{V}=0\}$. Since each term in (36) is non-negative, $\dot{V}=0$ implies
@@
The path-parameter rate $\dot{s}_{i}$ equals the component of velocity tangent to the path, that is, $\dot{s}_{i}=v_{i}\cos(\psi_{i}-\chi_{i}^{p})$, which reduces to $v_{i}$ when $\mathbf{x}_{i}\in\mathcal{P}_{m}\;(m=\{1,2\})$. Therefore, the asymptotic convergence to the path established in Section 3 justifies the assumption $\dot{s}_{i}=v_{i}$ used in Proposition 2.
@@ 5 Simulation Results
In this section, the proposed method is validated through numerical simulations in MATLAB for straight-path and sinusoidal-curve following. The initial UAV positions are uniformly sampled within a rectangular region bounded by the lines y=20 m,y=20 m,x=20 m,x=20 $\text{ m},y=20\text${ $\text{ m},x=-20\text${ m}. The safety distance of the UAVs and activation radius in (18) are considered to be $d_{\text{safe}}=0.4$ m and $R_{s}=1.5$ m, respectively. The nominal speed $v_{\text{nom}}=3$ m/s and $\kappa=1$ m/s, and the gain values $k_{r}=11,k_{g}=0.05,k_{\psi}=2.3$. The desired uniform spacing $d_{eq}$ is considered as 4 m. Further, we define the inter-UAV distance parameter $\mathcal{E}$ as the minimum distance between any UAV pair at a given time, that is, $\mathcal{E}=\min_{i\neq j}d_{i,j}$.
Figure 2: UAV trajectories for straight line path following (Ui represents the plot for the $i$th UAV).
Figure 3: UAV control input and spacing error for straight line path following: angular velocity input (top), linear velocity (middle), and spacing error (bottom) .
Figure 4: Inter-UAV distance parameter $\mathcal{E}$ for straight line path and sinusoidal curve following.
For the first scenario, straight line path-following is considered $(\mathcal{P}_{1}:x=0)$. The UAV trajectories are depicted in Fig. 2, which shows UAV positions at $t=0,20$ s and $40$ s. Fig. 3 shows the angular and linear velocity profiles of different UAVs, and the spacing error between consecutive UAVs. The angular velocity for all UAVs converges to 0, while the linear velocity converges to $v_{\text{nom}}=3$ m/s, while remaining within the bounds $(v_{\text{nom}}-\kappa,v_{\text{nom}}+\kappa)=(2,4)$ m/s. Since $\mathcal{E}>d_{\mathrm{safe}}=0.4$ m at all times, as shown in Fig. 4, no collision occurs between UAVs.
@@ 6 Conclusion
This paper presents a collision avoidance strategy for multiple UAVs following an open reference path while maintaining equal spacing along it. Two types of reference paths, straight line and sinusoidal curve, are considered, and a vector-field guidance law based on an arcsine function of the cross-track error drives each UAV to its reference path. Collision avoidance is achieved using a rotational repulsion mechanism that drives each UAV away from its neighbors based on relative bearing and proximity. The uniform spacing between UAVs along the path is maintained using a linear velocity control law that regulates each UAV’s speed based on the spacing error. Numerical simulations with 15 UAVs are performed to show the effectiveness of the proposed method. Future work could explore path-following for more complex open and closed paths. Benchmarking against existing potential function-based methods remains another direction for exploration.